OpenStudy (anonymous):

x^2 + 7x = –3 using quadratic formula

OpenStudy (anonymous):

x=1/2 (-7-sqroot37=-6.54 x=1/2(sqroot 37-7)=-0.45

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + 7x = –3\] \[\Large x^2 + 7x + 3 = 0\] \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(7)\pm\sqrt{(7)^2-4(1)(3)}}{2(1)}\] \[\Large x = \frac{-7\pm\sqrt{49-(12)}}{2}\] \[\Large x = \frac{-7\pm\sqrt{37}}{2}\] \[\Large x = \frac{-7+\sqrt{37}}{2} \ \text{or} \ x = \frac{-7-\sqrt{37}}{2}\]